OpenStudy (anonymous):

Sam decides to build a square garden. If the area of the garden is 4x2 − 20x + 25 square feet, what is the length of one side of the garden?

OpenStudy (anonymous):

\[4x ^{2}-20x+25= 4x ^{2}-10x-10x+25=2x(2x-5)-5(2x-5)=(2x-5)(2x-5)\]

OpenStudy (anonymous):

=(2x-5)(2x-5) =(2x-5)^2 so x= side of square= 5/2 = 2.5